# Categorical Flow Matching on Statistical Manifolds

### v1

@@ 4.1 Toy Example: Swiss Roll on Simplex
Figure 3: Generated samples of the Swill roll on simplex dataset and NLL (lower is better).
We note that most existing models assume the target data to be delta measures (one-hot distributions) and cannot be applied to this task. Both DDSM [7] and DirichletFM [56] imposed strong prior assumptions on the data distribution as Dirichlet distributions, which failed to capture the complex geometric shape of the Swill roll data points as demonstrated in the generated samples. On the contrary, directly built upon the geometry of the statistical manifold, our SFM model successfully captured the detailed data geometry. As no data resides on the boundary, exact NLL can be obtained using Eq.13 and was averaged over all points in the dataset. Though linear flow matching could also capture the geometry of the data, our SFM outperformed it in terms of NLL.
@@ 4.3 Text8
The Text8 dataset [39] is a media-size character-level corpus consisting of a small vocabulary of 27, which includes 26 lowercase letters and the whitespace token. We followed the convention in previous work [6, 23] to use random chunks of length 256 in both training and evaluation without any preprocessing. We used a 12-layer transformer-based predictor modified from [47] with additional time embeddings. As our exact likelihood is not directly comparable to bits-per-character (BPC) reported in previous work, we additionally calculated the pseudo-BPC inspired by [50]. See Appendix B.2 for additional details. All generated samples and NLL were estimated using 256 Euler steps. Quantitative results of BPC and NLL are provided in Tab.3 and generated texts are provided in Appendix D.2. The results for the transformer-based autoregressive language models are also provided as a reference. Our proposed SFM achieved the second-best performance among all diffusion or flow baselines. We also noted a significant performance gap between SFM and the naive linear flow matching on simplex, which highlighted the importance of capturing the intrinsic geometric properties of the underlying statistical manifold. Optimal transport does not significantly affect the final performance here, which we presume might be due to the long training stage in which vector fields were well-explored.
Table 2: BPC and NLL of different models on Text8. Results marked * are taken from corresponding papers.
@@
Table 3: SP-MSE on the generated promoter DNA sequences. * are from [7] and are from [56].
@@ 5 Related Work
As we put a special interest in discrete generation, we start with the existing work on discrete generation. We first list a few traditional autoregressive models [17, 47, 59] and masked language models [19, 50] but will skip them as we mainly focus on the diffusion and flow matching models. Existing diffusion and flow-based discrete generative models can be categorized into two main groups. The first group relies on stochastic jumps of Markov chains by treating discrete classes as Markov states. By choosing a proper transition matrix, the target one-hot distribution can be gradually noised into the stationary distribution. Noticeably, this approach can also adopt an additional absorbing state to mimic the mask token in masked language modeling [6]. D3PM [6] adapted the discrete-time Markov chain with the diffusion framework and SEDD [37] proposed a more efficient training scheme by learning the score entropy between different states. [12, 51] extended it to the continuous-time Markov chain by using the infinitesimal generator (rate matrix) instead, and [13] further applied the flow matching framework. Although the transition matrix or the rate matrix determines the entire diffusion trajectory, there is no canonical way of choosing it for different generation tasks to control the mixing rate. Also, due to the presence of discrete jumps of the Markov chain, exact likelihood calculation is no longer feasible. Only variational bounds were derived in [6, 13]. The other group directly works with the probability simplex or the logit space with certain assumptions of the underlying geometric structure. As an example, our linear flow matching assumes a Euclidean geometry on the probability simplex and is often used as a baseline in previous work [56]. The multinomial diffusion [27] assumed a Euclidean geometry on the logits space so interpolation became multiplication back on the probability simplex. Dirichlet diffusion (DDSM) [7] and Dirichlet flow matching (DirichletFM) [56] exerted a strong prior assumption on the probability path as the Jacobi diffusion process. We provide a more detailed analysis of these models in Appendix A.4. For models that directly operate on the logit space, the Bayesian flow network (BFN) assumed Gaussian distributions on the logit space with Bayesian update rules. [38] also assumed a Euclidean geometry in the logit space with targets as soft one-hot logits. The assumptions made in these models may not always hold, e.g., for the Swill roll on simplex dataset in Fig.3. These assumptions also did not necessarily capture the true geometry of the underlying statistical manifold. In contrast, our proposed SFM framework does not exert any strong prior assumptions and is aware of the intrinsic geometry by following the geodesics.
@@ A.1 Riemannian Manifold
A Riemannian manifold $\mathcal{M}$ is a real, smooth manifold equipped with a positive definite inner product $g$ on the tangent space $T_{x}(\mathcal{M})$ at each point $x\in\mathcal{M}$. Let $T\mathcal{M}=\bigcup_{x\in\mathcal{M}}T_{x}(\mathcal{M})$ be the tangent bundle of the manifold $\mathcal{M}$, a time-dependent vector field on $\mathcal{M}$ is a mapping $u_{t}:[0,1]\times\mathcal{M}\to T\mathcal{M}$ where $u_{t}(x)\in T_{x}(\mathcal{M})$. A geodesic is a locally distance-minimizing curve on the manifold. The existence and the uniqueness of the geodesic state that for any point $x\in\mathcal{M}$ and for any tangent vector $u\in T_{x}(\mathcal{M})$, there exists a unique geodesic $\gamma:[0,1]\to\mathcal{M}$ such that $\gamma(0)=x$ and $\gamma^{\prime}(0)=u$. The exponential map $\exp:\mathcal{M}\times T\mathcal{M}\to\mathcal{M}$ is uniquely defined to be $\exp_{x}(u):=\gamma(1)$. The logarithm map $\log:\mathcal{M}\times\mathcal{M}\to T\mathcal{M}$ is defined as the inverse mapping of the exponential map such that $\exp_{x}(\log_{x}(y))\equiv y,\forall x,y\in\mathcal{M}$ [^1] Note that the formulation here is slightly different from Sec.2.1 where we fix a point $\mu$ on $\mathcal{P}$.. With the exponential map and logarithm map, the time-dependent flow can be compactly written as time interpolation along the geodesic:
@@
It can be demonstrated that the above flow indeed traces the geodesic between $x_{0},x_{1}$ with linearly decreasing geodesic distance $d_{g}(x_{t},x_{1})=(1-t)d_{g}(x_{0},x_{1})$ [15]. For an $n$-dimensional Riemannian manifold, the geodesic can be obtained by solving the geodesic equation
$\sum_{j=1}^{n}g_{ji}\frac{d^{2}x_{j}}{dt^{2}}+\sum_{j=1}^{n}\sum_{k=1}^{n}\Gamma_{ijk}\frac{dx_{j}}{dt}\frac{dx_{k}}{dt}=0,\quad i=1,2,\dots,n$ (16)
where $x_{i}$ are local coordinates of the geodesic and $\Gamma_{ijk}$ are the Christoffel symbols of the first kind defined by
$\Gamma_{ijk}=\frac{1}{2}\left(\frac{\partial g_{ji}}{\partial x_{k}}+\frac{\partial g_{ji}}{\partial x_{j}}-\frac{\partial g_{jk}}{\partial x_{i}}\right),\quad i,j,k=1,2,\dots,n$ (17)
The closed-form expressions for the geodesic are generally not available. [42] curated a wide range of common statistical manifolds for parameterized families of both the discrete and the continuous distributions. We will focus on the statistical manifold of categorical distributions and also the spherical manifold, as the latter is closely related to the former via the diffeomorphism $\pi$ in Eq.(5).
@@ A.2 Spherical Manifold
The tangent space $T_{x}(S^{n-1}_{+})=\{u|\langle u,x\rangle=0\}$ is a $(n-1)$-dimensional hyperplane perpendicular to the vector $x$. The geodesic on the sphere follows the great circle between two points, and the geodesic distance can be calculated in Eq.7. The corresponding exponential map can be calculated as:
@@ A.3 Statistical Manifold of Categorical Distributions
The statistical manifold of categorical distributions $\mathcal{P}(\mathcal{X})$ is closely related to $S^{n-1}_{+}$ by the diffeomorphism $\pi$ in Eq.(5). In fact, $\pi$ is an isometry between the Fisher information metric on $\mathcal{P}$ and the standard induced metric on $S^{n-1}_{+}$ up to a constant scaling factor of $1/4$. To see this, we have the following proposition:
@@
We also noted the relation between the Fisher information metric defined in Eq.1 and the canonical inner product defined in Eq.4 for general statistical manifolds. Let $\mathcal{M}$ be an $n$-dimensional differentiable manifold and consider an embedding defined by $\mathcal{M}\hookrightarrow\mathcal{P(M)},\theta\mapsto p(\theta)=\sum_{i\in\mathcal{X}}p_{i}(\theta)\delta^{i}$. The pullback of the Fisher metric $g$ defines a metric on $\mathcal{M}$. For $u,v\in T_{\theta}(\mathcal{M})$, we have
@@
which gives the more common definition of the Fisher information matrix in Eq.1. This relation also generally holds for a continuous sample space $\mathcal{X}$ but requires more careful handling with measure theory tools. We refer interested readers to mathematical references, e.g., Chapter 3.1 in [8].
@@ A.4 Other Statistical Manifolds
  • DDSM and DirichletFM have strong prior assumptions that the categorical distributions follow some specific forward noising process which always terminates at (or near) one-hot distributions. This assumption generally holds for discrete generation, but cannot be generalized to more complex geometry on the probability simplex, as we have demonstrated in our toy example of the Swill roll on simplex dataset in Fig.3.
@@
Therefore, our model can also be interpreted as multinomial flow matching on the corresponding statistical manifold. The previous work of multinomial diffusion (argmax flow) [27] transformed the one-hot distribution to $ne_{k}-1$ in the logit space (soft one-hot logits) and assumed the common Euclidean structure of the logit space for constructing the diffusion process. This assumption did not consider the intrinsic geometry of the underlying statistical manifold and also had the inaccessibility issue of the delta measure as DirichletFM. We have demonstrated that SFM consistently outperformed the naive multinomial diffusion on the Text8 datasets, indicating the effectiveness of tracing the true geodesics.
@@ Appendix B Likelihood Calculation
In this section, we provide more concrete mathematical details regarding the exact likelihood calculation in the main text. We first make a distinction between negative log-likelihood (NLL) and bits-per-dimension (BPD, including bits-per-character, BPC) in this work. As we are dealing with probability distributions over the statistical manifold of probability measures, it is important for us to keep the difference between the probability spaces $\mathcal{P}(\mathcal{P}(\mathcal{X}))$ and $\mathcal{P}(\mathcal{X})$. In our notation, we use $\mu,\nu$ to denote elements in $\mathcal{P}(\mathcal{X})$, or categorical distributions, and use $p,q$ to denote elements in $\mathcal{P}(\mathcal{P}(\mathcal{X}))$, or distributions over categorical distributions. We further use NLL to refer to $-\log p(\mu)$ for a specific $\mu$ and BPD for categorical likelihood $-\log_{2}p(\delta|\mu)$ where $\delta$ is the ground truth delta measure (one-hot distribution). Note this definition of BPD (BPC) follows the convention in autoregressive language models where the conditional probabilities for the next token are calculated and averaged [17, 22]. In contrast, most previous work drew equivalence between these two concepts.
@@ B.1 Negative Log-Likelihood (NLL) Calculation
It is easy to see the density approaches infinity as $\theta$ approaches 0 or 1. Note that the arcsine distribution is a special case of beta distribution $\mathrm{Beta}(\frac{1}{2},\frac{1}{2})$, which is in turn a special case of Dirichlet distribution $\mathrm{Dir}(\frac{1}{2},\frac{1}{2})$. Each $\theta$ can be naturally viewed as the parameter for the Bernoulli distribution (2-class categorical distribution), so the arcsine distribution indeed defines a distribution over Bernoulli distributions with negative NLLs near the boundary. In fact, for discrete data, the target data always come in as delta measures, which indicates that the target distribution must be a convex combination of delta measures with infinite likelihood: $q=\sum_{i=1}^{n}\theta_{i}\delta_{i}$ where $\delta_{i}:=\delta_{\delta^{i}}$ is the delta measure over the underlying delta measure $\delta^{i}$ on the discrete class $i$.
@@
We followed [7] to derive a variational bound as marginalized over a small neighborhood of the delta measure in Eq.14 to obtain finite NLLs. The variational bound can be derived via the standard variational approach as
@@
The posterior probability $q(\mu|\delta)$ serves as the forward diffusion process in [6, 56] with a closed-form expression that depends on the timestep. In our Riemannian flow matching formulation, however, such a forward process is implicitly defined via the time-interpolation along the geodesics. Nonetheless, we do know that the conditional distribution should approach the delta measure when $t\to 1$ as the geodesic distance approaches 0: $p_{1}(\mu|\delta)\approx\delta$. Therefore, as we are marginalized over a small neighborhood, we can define a time-dependent posterior distribution $q_{t}(\mu|\delta)$ that approaches $\delta$ as $t\to 1$. In principle, any forward probability that approaches the delta measure at $t=1$ with an analytical likelihood can be used to derive the lower bound. For example, we can follow [7] to use the Jacobi diffusion process. To avoid the need for expensive Jacobi polynomial calculation, we instead use simple indicator measures over the simplex as
@@
where $p_{0}$ is the base probability of the uniform distribution over the simplex, $\mu_{i}=\theta_{i}$ is the $i$-th component of the categorical distribution of $\mu$. In other words, $q_{t}(\mu|\delta^{k})$ is the time-interpolated probability between the delta measure and the uniform distribution: $q_{t}(\mu|\delta^{k})=t\delta_{k}+(1-t)U\Delta^{n-1}$. We have
@@
The diffeomorphism $\pi$ in Eq.5 induces a push-forward measure on $S^{n-1}_{+}$ which can be described by the change of measure identity
@@
The model likelihood can be viewed as the push-forward measure of the learned flow:
@@
For manifolds that can be embedded in the ambient Euclidean space (e.g., simplex and sphere), the Riemannian divergence can be calculated as the normal Euclidean divergence $\mathrm{div}_{g}=\mathrm{div}$ [15]. We further use Hutchinson’s trace estimator [29] to efficiently estimate the divergence as
@@
The expectation can be efficiently computed using the vector-Jacobian product.
Combining Eq.(30), (31), (33), (37), (38) with Eq.(14) and (13), we can efficiently calculate the NLL given arbitrary delta measures as input. The NLL calculation scheme is visualized in Fig.2 and described in Alg.1.
@@
2:Apply the diffeomorphism in Eq.5 to obtain $\tilde{x}_{1}=\pi(\tilde{\mu}_{1})$ and calculate $\log|\det d\pi^{-1}(\tilde{x}_{1})|$.

### v2

@@ 4.1 Toy Example: Swiss Roll on Simplex
Figure 3: Generated samples of the Swiss roll on simplex dataset and NLL (lower is better). The NLLs are estimated using Hutchinson’s trace estimator, whereas those in the parenthesis are exact.
We note that most existing models assume the target data to be Dirac measures (one-hot distributions) and cannot be applied to this task. Both DDSM [7] and DirichletFM [59] imposed strong prior assumptions on the data distribution as Dirichlet distributions, which failed to capture the complex geometric shape of the Swiss roll data points as demonstrated in the generated samples. On the contrary, directly built upon the geometry of the statistical manifold, our SFM model successfully captured the detailed data geometry. As no data resides on the boundary, exact NLL can be obtained using Eq.(13) and was averaged over all points in the dataset. Though linear flow matching could also capture the geometry of the data, our SFM outperformed it in terms of NLL.
@@ 4.3 Text8
The Text8 dataset [40] is a medium-size character-level corpus consisting of a small vocabulary of 27, which includes the 26 lowercase letters and the whitespace token. We followed the convention in previous work [6, 23] to use random chunks of length 256 in both training and evaluation without any preprocessing. We used a 12-layer diffusion transformer (DiT) [48] based predictor, similar to the one used in [38]. As our exact likelihood is not directly comparable to bits-per-character (BPC) reported in previous work which relies on an alternative variational bound for the likelihood, we additionally calculate such BPC inspired by [51]. See Appendix B.2 for additional details. All generated samples and NLL were estimated using the Dopri5 ODE solver. Quantitative results of BPC are provided in Tab.3 and generated texts are provided in Appendix D.3. The results for the autoregressive language models are also provided as a reference (Discrete Flow [63] is based on autoregressive normalizing flow).
Note that, unlike all of the other baselines, SFM and LinearFM do not directly optimize such a BPC as a training objective, so such an evaluation metric is unfavorable for our model (see Appendix B.2). Despite such an unfavorable evaluation, our proposed SFM still achieved the second-best performance among other diffusion and flow baselines that were directly trained with cross-entropy losses. We also noted a significant performance gap between SFM and the naive linear flow matching on simplex, highlighting the importance of capturing the intrinsic geometric properties of the underlying statistical manifold. Optimal transport does not significantly affect the final performance here, which we presume might be due to the long training stage in which vector fields were well-explored. Additional evaluation following [12] is provided in Appendix D.1, in which SFM achieved the best generation entropy as evaluated by the pre-trained GPT-J-6B model [66].
Table 2: BPC on Text8. Results marked * are taken from the corresponding papers.
@@
Table 3: SP-MSE (as evaluated by Sei [13]) on the generated promoter DNA sequences. Results marked * are from [7] and results marked are from [59].
@@ 5 Related Work
As we put a special interest in discrete generation, we start with the existing work on discrete generation. We first list a few traditional autoregressive models [16, 49, 63] and masked language models [18, 53] but will skip them as we mainly focus on the diffusion and flow matching models. Existing diffusion and flow-based discrete generative models can be categorized into two main groups. The first group relies on stochastic jumps of Markov chains by treating discrete classes as Markov states. By choosing a proper transition matrix, the target one-hot distribution can be gradually noised into the stationary distribution. Noticeably, this approach can also adopt an additional absorbing state to mimic the mask token in masked language modeling [6]. D3PM [6] adapted the discrete-time Markov chain with the diffusion framework and SEDD [38] proposed a more efficient training scheme by learning the score entropy between different states. [11, 54] extended it to the continuous-time Markov chain by using the infinitesimal generator (rate matrix) instead, and [12] further applied the flow matching framework. Although the transition matrix or the rate matrix determines the entire diffusion trajectory, there is no canonical way of choosing it for different generation tasks to control the mixing rate. Also, due to the presence of discrete jumps of the Markov chain, exact likelihood calculation is no longer feasible. Only variational bounds were derived in [6, 12]. The other group directly works with the probability simplex or the logit space with certain assumptions of the underlying geometric structure [35]. As an example, our linear flow matching assumes a Euclidean geometry on the probability simplex and is often used as a baseline in previous work [59]. The multinomial diffusion [27] assumed a Euclidean geometry on the logits space so interpolation became multiplication back on the probability simplex. Dirichlet diffusion (DDSM) [7] and Dirichlet flow matching (DirichletFM) [59] exerted a strong prior assumption on the probability path as the Jacobi diffusion process. We provide a more detailed analysis of these models in Appendix A.4. For models that directly operate on the logit space, the Bayesian flow network (BFN) assumed Gaussian distributions on the logit space with Bayesian update rules. [39] also assumed a Euclidean geometry in the logit space with targets as soft one-hot logits. The assumptions made in these models may not always hold, e.g., for the Swiss roll on simplex dataset in Fig.3. These assumptions also did not necessarily capture the true geometry of the underlying statistical manifold. In contrast, our proposed SFM framework does not exert any strong prior assumptions and is aware of the intrinsic geometry by following the geodesics.
@@ A.1 Riemannian Manifold
A Riemannian manifold $\mathcal{M}$ is a real, smooth manifold equipped with a positive definite inner product $g$ on the tangent space $T_{x}(\mathcal{M})$ at each point $x\in\mathcal{M}$. Let $T\mathcal{M}=\bigcup_{x\in\mathcal{M}}T_{x}(\mathcal{M})$ be the tangent bundle of the manifold $\mathcal{M}$, a time-dependent vector field on $\mathcal{M}$ is a mapping $u_{t}:[0,1]\times\mathcal{M}\to T\mathcal{M}$ where $u_{t}(x)\in T_{x}(\mathcal{M})$. A geodesic is a locally distance-minimizing curve on the manifold. The existence and the uniqueness of the geodesic state that for any point $x\in\mathcal{M}$ and for any tangent vector $u\in T_{x}(\mathcal{M})$, there exists a unique geodesic $\gamma:[0,1]\to\mathcal{M}$ such that $\gamma(0)=x$ and $\gamma^{\prime}(0)=u$. The exponential map $\exp:\mathcal{M}\times T\mathcal{M}\to\mathcal{M}$ is uniquely defined to be $\exp_{x}(u):=\gamma(1)$. The logarithm map $\log:\mathcal{M}\times\mathcal{M}\to T\mathcal{M}$ is defined as the inverse mapping of the exponential map such that $\exp_{x}(\log_{x}(y))\equiv y,\forall x,y\in\mathcal{M}$***Note that the formulation here is slightly different from Sec.2.1 where we fix a point $\mu$ on $\mathcal{P}$.. With the exponential map and logarithm map, the time-dependent flow can be compactly written as time interpolation along the geodesic:
@@
It can be demonstrated that the above flow indeed traces the geodesic between $x_{0},x_{1}$ with linearly decreasing geodesic distance $d_{g}(x_{t},x_{1})=(1-t)d_{g}(x_{0},x_{1})$ [14]. For an $n$-dimensional Riemannian manifold, the geodesic can be obtained by solving the geodesic equation (using the Einstein summation convention):
$\frac{d^{2}x^{k}}{dt^{2}}+\Gamma_{ij}^{k}\frac{dx^{i}}{dt}\frac{dx^{j}}{dt}=0,\quad k=1,2,\dots,n$ (16)
where $x^{i}$ are local coordinates of the geodesic and $\Gamma_{ij}^{k}$ are the Christoffel symbols defined by
$\Gamma_{ij}^{k}=\frac{1}{2}g^{km}\left(\frac{\partial g_{mi}}{\partial x^{j}}+\frac{\partial g_{mj}}{\partial x^{i}}-\frac{\partial g_{ij}}{\partial x^{m}}\right),\quad i,j,k=1,2,\dots,n$ (17)
where $g^{ij}$ is the inverse metric such that $g^{ij}g_{jk}=\delta_{k}^{i}$. The closed-form expressions for the geodesic are generally not available. [43] curated a wide range of common statistical manifolds for parameterized families of both the discrete and the continuous distributions. We will focus on the statistical manifold of categorical distributions and also the spherical manifold, as the latter is closely related to the former via the diffeomorphism $\pi$ in Eq.(5).
@@ A.2 Spherical Manifold
The tangent space $T_{x}(S^{n-1}_{+})=\{u|\langle u,x\rangle=0\}$ is a $(n-1)$-dimensional hyperplane perpendicular to the vector $x$. The geodesic on the sphere follows the great circle between two points, and the geodesic distance can be calculated in Eq.(7). The corresponding exponential map can be calculated as:
@@ A.3 Statistical Manifold of Categorical Distributions
By calculating the Fisher information matrix for a categorical distribution, we can obtain the explicit form of the metric tensor as [43]
$g_{jk}(\mu)=\frac{\delta_{jk}}{\mu_{j}}+\frac{1}{\mu_{n}},\quad 1\leq i,j,\leq n-1$ (21)
where $\delta_{jk}$ is the Kronecker delta. Substituting this into $\langle u,v\rangle_{\mu}$ leads to the Riemannian inner product in Eq.(4). The statistical manifold of categorical distributions $\mathcal{P}(\mathcal{X})$ is closely related to $S^{n-1}_{+}$ by the diffeomorphism $\pi$ in Eq.(5). In fact, $\pi$ is an isometry between the Fisher information metric on $\mathcal{P}$ and the standard induced metric on $S^{n-1}_{+}$ up to a constant scaling factor of $1/4$. To see this, we have the following proposition:
@@
We specifically note the property that the Riemannian structure induced by the Fisher information metric leads to vector fields more parallel to the boundaries. This can also be derived mathematically from the logarithm map in Eq.(24). Consider the direction term $\sqrt{\mu\odot\nu}-\langle\sqrt{\mu},\sqrt{\nu}\rangle\mu$. For $\mu$ close to the boundary with $\mu_{k}\approx 0$, its corresponding vector field will also have a close to $u_{k}\approx 0$ component, which is different from linear flow matching’s fixed $\nu-\mu$. We hypothesize that one potential benefit of such a curved geometry over the naive Euclidean geometry is that the former helps prevent overshooting across the boundaries. Specifically, consider a target point on the boundary. The Euclidean vector field will continue to push the points outside the manifold, whereas the Riemannian vector field tends to travel parallel to that boundary to prevent going across the boundary. Once overshooting happens during sampling, the model may exhibit undefined behavior as it was never trained on the points outside the manifold.
We also noted the relation between the Fisher information metric defined in Eq.(1) and the canonical inner product defined in Eq.(4) for general statistical manifolds. Let $\mathcal{M}$ be an $n$-dimensional differentiable manifold and consider an embedding defined by $\mathcal{M}\hookrightarrow\mathcal{P(M)},\theta\mapsto p(\theta)=\sum_{i\in\mathcal{X}}p_{i}(\theta)\delta^{i}$. The pullback of the Fisher metric $g$ defines a metric on $\mathcal{M}$. For $u,v\in T_{\theta}(\mathcal{M})$, we have
@@
which gives the more common definition of the Fisher information matrix in Eq.(1). This relation also generally holds for a continuous sample space $\mathcal{X}$ but requires more careful handling with measure theory tools. We refer interested readers to mathematical references, e.g., Chapter 3.1 in [8].
@@ A.4 Other Statistical Manifolds
  • DDSM and DirichletFM have strong prior assumptions that the categorical distributions follow some specific forward noising process which always terminates at (or near) one-hot distributions. This assumption generally holds for discrete generation, but cannot be generalized to more complex geometry on the probability simplex, as we have demonstrated in our toy example of the Swiss roll on simplex dataset in Fig.3.
@@
Therefore, our model can also be interpreted as multinomial flow matching on the corresponding statistical manifold. The previous work of multinomial diffusion (argmax flow) [27] transformed the one-hot distribution to $ne_{k}-1$ in the logit space (soft one-hot logits) and assumed the common Euclidean structure of the logit space for constructing the diffusion process. This assumption did not consider the intrinsic geometry of the underlying statistical manifold and also had the inaccessibility issue of the Dirac measure as DirichletFM. We have demonstrated that SFM consistently outperformed the naive multinomial diffusion on the Text8 datasets, indicating the effectiveness of tracing the true geodesics.
@@ Appendix B Likelihood Calculation
In this section, we provide more concrete mathematical details regarding the exact likelihood calculation in the main text. We first make a distinction between negative log-likelihood (NLL) and bits-per-dimension (BPD, including bits-per-character, BPC) in this work. As we are dealing with probability distributions over the statistical manifold of probability measures, it is important for us to keep the difference between the probability spaces $\mathcal{P}(\mathcal{P}(\mathcal{X}))$ and $\mathcal{P}(\mathcal{X})$. In our notation, we use $\mu,\nu$ to denote elements in $\mathcal{P}(\mathcal{X})$, or categorical distributions, and use $p,q$ to denote elements in $\mathcal{P}(\mathcal{P}(\mathcal{X}))$, or distributions over categorical distributions. We further use NLL to refer to $-\log p(\mu)$ for a specific $\mu$ and BPD for categorical likelihood $-\log_{2}p(\delta|\mu)$ where $\delta$ is the ground truth Dirac measure (one-hot distribution). Note this definition of BPD (BPC) follows the convention in autoregressive language models where the conditional probabilities for the next token are calculated and averaged [16, 22]. In contrast, most previous work drew equivalence between these two concepts.
@@ B.1 Negative Log-Likelihood (NLL) Calculation
It is easy to see the density approaches infinity as $\theta$ approaches 0 or 1. Note that the arcsine distribution is a special case of beta distribution $\mathrm{Beta}(\frac{1}{2},\frac{1}{2})$, which is in turn a special case of Dirichlet distribution $\mathrm{Dir}(\frac{1}{2},\frac{1}{2})$. Each $\theta$ can be naturally viewed as the parameter for the Bernoulli distribution (2-class categorical distribution), so the arcsine distribution indeed defines a distribution over Bernoulli distributions with negative NLLs near the boundary. In fact, for discrete data, the target data always come in as Dirac measures, which indicates that the target distribution must be a convex combination of Dirac measures with infinite likelihood: $q=\sum_{i=1}^{n}\theta_{i}\delta_{i}$ where $\delta_{i}:=\delta_{\delta^{i}}$ is the Dirac measure over the underlying Dirac measure $\delta^{i}$ on the discrete class $i$.
@@
We followed [7] to derive a variational bound as marginalized over a small neighborhood of the Dirac measure in Eq.(14) to obtain finite NLLs. The variational bound can be derived via the standard variational approach as
@@
The posterior probability $q(\mu|\delta)$ serves as the forward diffusion process in [6, 59] with a closed-form expression that depends on the timestep. In our Riemannian flow matching formulation, however, such a forward process is implicitly defined via the time-interpolation along the geodesics. Nonetheless, we do know that the conditional distribution should approach the Dirac measure when $t\to 1$ as the geodesic distance approaches 0: $p_{1}(\mu|\delta)\approx\delta$. Therefore, as we are marginalized over a small neighborhood, we can define a time-dependent posterior distribution $q_{t}(\mu|\delta)$ that approaches $\delta$ as $t\to 1$. In principle, any forward probability that approaches the Dirac measure at $t=1$ with an analytical likelihood can be used to derive the lower bound. For example, we can follow [7] to use the Jacobi diffusion process. To avoid the need for expensive Jacobi polynomial calculation, we instead use simple indicator measures over the simplex as
@@
where $p_{0}$ is the base probability of the uniform distribution over the simplex, $\mu_{i}=\theta_{i}$ is the $i$-th component of the categorical distribution of $\mu$. In other words, $q_{t}(\mu|\delta^{k})$ is the time-interpolated probability between the Dirac measure and the uniform distribution: $q_{t}(\mu|\delta^{k})=t\delta_{k}+(1-t)U\Delta^{n-1}$. We have
@@
The diffeomorphism $\pi$ in Eq.(5) induces a pushforward measure on $S^{n-1}_{+}$ which can be described by the change of measure identity
@@
The model likelihood can be viewed as the pushforward measure of the learned flow:
@@
For manifolds that can be embedded in the ambient Euclidean space (e.g., simplex and sphere), the Riemannian divergence can be calculated as the normal Euclidean divergence $\operatorname{div}_{g}=\operatorname{div}$ [14]. We further use Hutchinson’s trace estimator [29] to efficiently estimate the divergence as
@@
The expectation can be efficiently computed using the vector-Jacobian product. For data with a unitary dimension $D=1$ (e.g., Swiss roll on simplex), we also calculate the exact divergence using the equality $\operatorname{div}v=\operatorname{Tr}(J(v))$, where $J(v)$ is the Jacobian of the vector field $v$. The computational cost for the exact divergence calculation scales quadratically with the data dimension $D$, as the interdependence between data dimensions makes it computationally expensive to calculate the Jacobian. Therefore, we only demonstrated the NLL with exact divergence calculation for the Swiss roll dataset in Fig.3, in which Hutchinson’s trace estimator could make a decent estimation of the exact divergence.
Combining Eq.(31), (32), (34), (38), and (39) with Eq.(14) and (13), we can efficiently calculate the NLL given arbitrary Dirac measures as input. The NLL calculation scheme is visualized in Fig.2 and described in Alg.1.
@@
2:Apply the diffeomorphism in Eq.(5) to obtain $\tilde{x}_{1}=\pi(\tilde{\mu}_{1})$ and calculate $\log|\det d\pi^{-1}(\tilde{x}_{1})|$.